OpenStudy (anonymous):

Write the metric unit that makes the statement true. 2. 5,941 mm = 5.941 __

OpenStudy (anonymous):

Well, it's looking for another metric unit, so you know it'll be a conversion factor of \(10^x\). You need to know what \(x\) is. \[5941 \times \frac{1}{?} = 5.941\]

OpenStudy (anonymous):

Or, put another way: \[5941 \times \frac{1}{10^{x}} = 5.941\] Solve for \(x\).

OpenStudy (anonymous):

m km cm dm these are the possible answers

OpenStudy (anonymous):

You need to determine the order of magnitude difference between 5941 and 5.491. Then, you can easily discern which units you've converted to.

OpenStudy (anonymous):

do you know the answer?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

what is it??

OpenStudy (anonymous):

How will telling you help? Then you still won't know how to do the problem.

OpenStudy (anonymous):

grr

OpenStudy (anonymous):

I've given you the steps.

OpenStudy (anonymous):

Make it easier on yourself. Round up to 6000. 6000 divided by what equals 6.000?

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

Okay... 6000 divided by what equals 600?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Good. And 6000 divided by what is 60? From that, 6000 divided by what is 6?

OpenStudy (anonymous):

100 then 1000

OpenStudy (anonymous):

Good. So, the conversion factor from 5941 to 5.941 is 1/1000, right?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

So, which metric unit is 1000 mm?

OpenStudy (anonymous):

\[5941 \times \frac{1 \text{ ??}}{1000 \text{ mm}} = 5.941 ??\]

OpenStudy (anonymous):

idk metric units

OpenStudy (anonymous):

Oh... Yeah, that's a bit important for this one. :)

OpenStudy (anonymous):

ya thats why i asked

OpenStudy (anonymous):

That's odd. I would imagine you would only be assigned a question like this if you were learning metric.

OpenStudy (anonymous):

me too